I have posted about this problem once and it was said to not be able to be solved but it is to be solved because it is a specific question I need help with. please and thank you if you can help me!!
Create a trail mix with three ingredients; peanuts, fruit, and cashews. The cost per pound of peanuts are $3, fruit $6, and cashews $3. One ingredient must have twice the weight of other two. (that part can be made up, I was not sure what to make those numbers). using systems of equations and or matrices how many pounds of each should be used to make a 140 lb mix at 6$ a pound?

Answer by TimothyLamb(4379)

x = pounds of peanuts
y = pounds of fruit
z = pounds of cashews
---
"one ingredient must have twice the weight of other two" ...
arbitrarily choose the weight of peanuts to be twice the sum of the weights of fruit and cashews:
x = 2(y + z)
---
sum of weights:
x + y + z = 140
---
sum of costs:
3x + 6y + 3z = 6*140
---
put the system of linear equations into standard form
---
x = 2y + 2z
x + y + z = 140
3x + 6y + 3z = 840
---
x - 2y - 2z = 0
x + y + z = 140
3x + 6y + 3z = 840
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = pounds of peanuts = 93.333333
y = pounds of fruit = 140
z = pounds of cashews = -93.333333
---
as you can see, pounds of cashews comes out negative, which is impossible in real life.
this means that the problem statement is fundamentally incorrect.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---